OpenStudy (anonymous):

Write the rational number as the ratio of two integers in simplest form. 0.395959595... detailed steps would be appreciated :)

OpenStudy (anonymous):

0.95959595... = 95/99 0.0959595... = 95/99*(1/10) = 95 / 990 0.3 = 3/10 0.3959595... = 0.3 + 0.0959595 = 3/10 + 95/990 = 196/495

OpenStudy (anonymous):

why is it over 99?

OpenStudy (anonymous):

Anytime you have 0.something and that something repeats you can put the pattern over 9s. The number of 9s you use depends on the length of the pattern. For example 0.44444 = 4/9 0.182182182... = 182/999 0.418534185341853... = 41853/99999

OpenStudy (nikvist):

\[x=0.39595959...\]\[10x=3.9595959...\]\[1000x=395.95959...\]\[1000x-10x=395.95959...-3.9595959...\]\[990x=392\]\[x=\frac{392}{990}=\frac{196}{495}\]

OpenStudy (anonymous):

oh, also why did you multiply it by 1/10?

OpenStudy (anonymous):

A lot of people use @nikvist 's method too. That works, too.

OpenStudy (anonymous):

Because we want 0.0959595...not 0.959595.... To move the decimal, times by 1/0 0.3959595...is 0.3 + 0.095959595

OpenStudy (anonymous):

oh okay now i understand thank you for all the help